OpenStudy (anonymous):

\[\int\limits_{}^{} dx /\sqrt{x ^{2}+2x+2} \log_{e} (1+x+\sqrt{x ^{2}+2x+2})\]

OpenStudy (anonymous):

\[\int\limits_{}^{} dx /\sqrt{x ^{2}+2x+2} \log_{e} (1+x+\sqrt{x ^{2}+2x+2})\]